OpenStudy (anonymous):

when serving in tennis a player tosses the tennis ball vertically in the . The height h of the ball after t seconds is given by the quadratic function h(t)=-5t^2+7t (the height is measured in meters from the point of the toss.) a. How high in the air does the ball go? b. Assume that the player hits the ball on its way down when it's 0.6m above the point of the toss. For how many seconds is the ball in the air between the toss and the serve?

OpenStudy (anonymous):

First sentence is in the air. Oops

OpenStudy (anonymous):

\[h(t)=-5t^2+7t\] max is the second coordinate of the vertex first coordinate of the vertex is always \(-\frac{b}{2a}\) which in your case is \[-\frac{7}{2\times (-5)}=\frac{7}{10}=.7\] second coordinate is what you get when you replace \(t\) by \(.7\) and that is your maximum height

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

What do I do next?

OpenStudy (anonymous):

pleaase hellppp